Journalist Report – January 12th

Crew 219 Journalist Report, 12 JAN 2020

Author: Nathan Hadland

Sol 7

A DONSA To Remember

Ah, the DONSA. The day where you take a moment to breathe and recharge
your batteries. Indeed, today allowed our minds and bodies to recover
from a week of productive science and maintenance tasks. The crew
slept in this morning and woke up to the sound of reggae as well as
the ever-present water pump. As I climbed down from the loft, I saw
our commander in a bright yellow Hawaiian shirt making a breakfast of
bacon and eggs. The rest of the crew emerged in their own uniquely
colored Hawaiian shirts and I chuckled at the irony of a group of
scientists on Mars hanging out as if they were laying on the beach
back in Florida. After sweeping the tunnels of snow and cleaning the
lower deck, I challenged Robinson (ASTRO) and Alejandro (ENG) to an
epic game of Super Smash Bros Ultimate. The battle was intense,
exhilarating, and nerve-racking. Robinson joked that we smacked him
off of Mars back to Florida. For lunch, Hannah (LSO) and Cynthia (GHO)
harvested some greens from the GreenHab and made a delicious fresh
salad along with sliders and fresh bread. For the rest of the day,
Robinson and Abdul (GEO) worked in the Science Dome on their samples.
I got to witness Robinson simulate a Martian dust storm to investigate
the most effective dust mitigation methods for optical mirrors in
astronomical observations. He will be presenting this work at a
conference later this spring! He also got some awesome images last
night for his astronomy research.

After working in the lab, I spent the rest of the afternoon reading my
book and listening to the ambient noise of Alejandro, Cynthia, and
Dave playing cards. We also got to enjoy some delicious cookies made
by Hannah and heavenly Chai tea made by Abdul. Later tonight, we will
be making a dinner of blueberry pancakes and hash browns. Tomorrow we
will be back to work and continuing sample collection. I am both
excited and humbled at the opportunity to be working in an environment
as magnificent as this and am eager to see what this week on the Red
Planet will reveal.

GreenHab Report – January 12th

Crew 219 GreenHab Report 12-JAN-20

Crew GreenHab Officer: Cynthia Montanez

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 24.3 °C; 18 %

Average Max: 28.3 °C, 19.3%

Average Min: 21.6 °C, 15.7%

07:45 A.M.

Floor Unit: 15 °C

Electronic: 24.1 °C

humidity 18%

Max: 26.2 °C; 20%

Min: 20.3 °C; 17%

3:10 P.M.

Floor Unit: 19 °C

Electronic: 24.5 °C

humidity 18%

Max: 29.4 °C; 19%

Min: 22.5 °C; 15%

5:54 P.M.

Floor Unit: 17 °C

Electronic: 24.4 °C

humidity 18%

Max: 29.4 °C; 19%

Min: 21.9 °C; 15%

Hours of supplemental light: Light system 7:00PM-12:00AM.

Daily water usage for crops: 4.59 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – 174.86 Gal.

Time(s) of watering for crops:

7:54 A.M.

SEASONAL PLANTING HISTORY:

Change to crops:

OTHER NOTES:

1st sprouts: N/A

Harvest: At 11:45 A.M., 74 grams of butter lettuce was harvested. At
12:07 P.M., 146 grams of swiss chard was harvested. At 12:11 P.M., 36
grams of purple carrots were harvested.

Other: N/A

Research Report – January 12th

Crew 219 Science Report 12-JAN-20
Crew Science Officer: Hannah Blackburn

1. Biometrics and Neurobehavioral Research

Nothing to report.

2. UAV

Nothing to report.

3. Dust Mitigation for Optical Mirrors

Simulated dust storm by manually applying regolith dust to mirrors.

4. Astrophotography of Celestial Bodies

See Sol 6 astronomy report and photos. Made color images of M32,
NGC1068, and NGC7318.

5. Remediation of Mars Regolith

Nothing to report.

6. Chemical and Mineralogical Composition of the MDRS Site

Samples were dried in the oven using glass beakers.

7. Protocols for the Discovery of Life on Mars

Nothing to report.

Glassware check out:

10 glass beakers (sizes between 200 and 500 ml) were used to dry
regolith samples.

A reminder to all crewmembers: There’s a $300 fine for using any
glassware material without MDRS Mission support permission.

Operations Report – January 12th

Operations Report

Crew 219 Operations Report 12-JAN-2020

Sol: 7

Name of person filing report: Alejandro Perez

Non-nominal systems: N/A

Generator: Run

Hours run: 17

From what time last night: 15:00

To what time this morning: 08:00

List of any additional daytime hours when the generator was run: N/A

Solar- SOC% (Before the generator is run at night): 82%

09:40 = 94%; 11:00 = 90%; 11:50 = 88%; 12:55 = 100%; 14:10 = 99%;
15:10 = 97%; 16:15 = 94%; 17:05 = 90%; 18:15 = 82%

Diesel Reading – 60%

Station Propane Reading- 40%

Ethanol Free Gasoline- N/A

Water (loft tank) (gal) – 50

Water Meter (units)- 0147603,1

Water (static tank) (gal) – 411

Static to loft Pump used –Yes

Water in Green Hab (gal): 174.86

Water in Science Dome (gal): 0

Toilet Tank Emptied –No

Diemos rover used: Still in the Work shop

Hours:

Beginning Charge:

End Charge:

Currently Charging:

Sojourner rover used: Yes, by Outpost

Hours:

Beginning Charge:100 %

End Charge: 21%

Currently Charging: Yes

Spirit Rover used- No

Starting Hours:

Beginning Charge:

Ending Hours:

Ending Charge:

Opportunity Rover used: Still in workshop

Hours:

Beginning Charge:

Ending Charge:

Currently Charging:

Curiosity Rover used: Yes, by Outpost

Starting Hours:

Beginning Charge: 100%

Ending Hours: 31%

Ending Charge: Yes

Notes on Rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use:

Oil Added? No

# hours the ATV’s were used today:

Notes on ATVs:

Habcar used and why, where?

CrewCar used and why, where?

General Notes and Comments:

Summary of the internet: Nominal

Summary of Suit and Radios: Nominal

Summary of Hab: SOS’s were delivered. Repair to loft tank feedline is
working properly, however, the joint connecting the pipe going into
the loft tank is still leaking even with the new Teflon tape added.

Summary of Science Dome operations: Temperatures are hovering around 7
°C during the day.

Summary of Ram operations: Nominal

Summary of any observatory issues: Nominal

Summary of Health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: ¾ in. M-M metal
threaded coupling for feedline pipe.

Sol Summary – January 12th

Crew 219 Sol Summary Report 12-01-2020

Sol: 7

Summary Title: Bacon, Burgers, Hawaiian Shirts and Mistaken Aces

Author’s name: Dave Masaitis

Mission Status: Nominal

Sol Activity Summary: Despite today being a DONSA (see yesterday’s Sol
Summary for definition), work did actually get done today. After a
breakfast of bacon and eggs, some crew members swept snow out of the
tunnels, while others swept and mopped mud from the Hab’s first floor.
A. Elnajdi performed some sample analysis in the Science Dome and R.
Raphael performed astronomic data reduction while H. Blackburn and C.
Montanez made bacon cheeseburgers and fresh salad with GreenHab greens
for lunch. Afterwards N. Hadland and R. Raphael simulated a dust storm
for our dust mitigation experiments, and others engaged in a lively
game of Rummy 500. C. Montanez took a victory, before the crew sat
down to finish reports and sip the world’s best chai tea, courtesy of
A. Elnajdi. The evening will hold a crew meeting to prepare for
tomorrow’s work, and then it’s back to business in the morning.

Look Ahead Plan: One EVA, continued sample analysis, and maintenance
if necessary

Anomalies in work: None

Weather: Cold and overcast most of the day, with some sunshine in the
late afternoon

Crew Physical Status: Nominal

EVA:

None

Reports to be filed:

Sol Summary

EVA Request #13

Operations Report

GreenHab Report

Research Report

Journalist’s Report

Astronomy Photos (ATTN: Dr. Zubrin)

Photos (Including Photo of the Day)

Support Requested:

– Refer to Operations Report